Find the probabilities on a standard normal curve P(z<0.45)


1
Expert's answer
2021-03-01T15:31:54-0500

"P(Z<0.45)=P(-\infty <Z<0)+P(0< Z<0.45)"

"=0.5+P(0<Z<0.45)"

"=0.5+0.1736" [ from standard normal distribution table]

"=0.6736"